# A Survey on Statistical Theory of Deep Learning: Approximation, Training Dynamics, and Generative Models

### v1

@@ 2.3 Non-parametric Classification via fully-connected networks
Lastly, a paper Hu et al. (2020) studied the convergence rate of excess risk of neural networks under the framework of Mammen and Tsybakov (1999). Specifically, the authors of the paper consider a Teacher-Student framework. In this setup, one neural network, called student net, is trained on data generated by another neural network, called teacher net. Adopting this framework can facilitate the understanding of how deep neural networks work as it provides an explicit target function with bounded complexity. Furthermore, assuming the target classifier to be a teacher network of an explicit architecture might provide insights into what specific architecture of the student classifier is needed to achieve an optimal excess risk. Under this setting, the rate of convergence is derived as $\tilde{\mathcal{O}}_{d}(n^{-\frac{2}{3}})$ for the excess risk of the empirical $0-1$ loss minimizer, given that the student network is deeper and larger than the teacher network. The authors use the notation $\tilde{\mathcal{O}}_{d}$ to denote the rate is dependent on input dimension $d$ in a logarithmic factor. When data are separable, the rate improves to $\tilde{\mathcal{O}}_{d}(n^{-1})$. In contrast, as it has already been shown in Mammen and Tsybakov (1999), under CAD assumption, the optimal rate is $\mathcal{O}(n^{-\alpha(q+1)/[\alpha(q+2)+(d-1)q]})$. Clearly, this rate suffers from “Curse of Dimensionality” but interestingly, coincides with the rate $\tilde{\mathcal{O}}_{d}(n^{-\frac{2}{3}})$ when $\alpha=1$ and $\beta\rightarrow{\infty}$. If we further allow $\alpha\rightarrow{\infty}$ (corresponding to separable data), the classical rate above recovers $\mathcal{O}(n^{-1})$. The paper also got the rate is un-improvable in a minimax sense by showing that the minimax lower bound has the same order as that of the upper bound (i.e., $\tilde{\mathcal{O}}_{d}(n^{-\frac{2}{3}})$).
3 Overparametrized neural networks
We consider the following shallow neural network $f_{\mathbf{W}}(\mathbf{x})$ with $M$ number of hidden neurons:
$f_{\mathbf{W}}(\mathbf{x})=\frac{\alpha}{M}\sum_{r=1}^{M}a_{r}\sigma(w_{r}^{\top}\mathbf{x}),$ (9)
where $\mathbf{x}\subset\mathcal{X}$ is an input vector, $\{w_{r}\}_{r=1}^{M}$ are the weights in the first hidden layer, and $\{a_{r}\}_{r=1}^{M}$ are the weights in the output layer. Let us denote the pair $\mathbf{W}:=\{(a_{r},w_{r})\}_{r=1}^{M}$. The network dynamic is scaled with the factor $\frac{\alpha}{M}$. If the network width (i.e., $M$) is small, the scaling factor has negligible effects on the network dynamics. But for the wide enough network (i.e., overparametrized setting), the scaling difference yields completely different behaviors in the dynamics. In this review, we will focus on two specific regimes: (1) Neural Tangent Kernel regime (Jacot et al., 2018; Du et al., 2019; Arora et al., 2019b) with $\alpha=\sqrt{M}$, (2) Mean Field regime (Nitanda and Suzuki, 2017; Chizat and Bach, 2018; Mei et al., 2019) with $\alpha=1$. Additionally, we will review a work that gives unifying views on these two regimes.
In this section, we focus on reviewing the papers that work on $\ell_{2}$-loss function :
$\mathcal{L}_{\mathbf{S}}\big{(}\mathbf{W}\big{)}=\frac{1}{2}\sum_{i=1}^{n}\big{(}y_{i}-f_{\mathbf{W}}(\mathbf{x}_{i})\big{)}^{2},$ (10)
and we update the model parameters $\mathbf{W}$ through the following gradient descent update rule. Let $\mathbf{W}_{0}$ be the initialized weight, and we consider the following gradient descent update rule with step-size $\eta>0$:
$\mathbf{W}_{(k)}=\mathbf{W}_{(k-1)}-\eta\nabla_{\mathbf{W}}\big{(}\mathcal{L}_{\mathcal{S}}(\mathbf{W}_{(k-1)})\big{)},\qquad k\geq 1.$ (11)
3.1 Neural Tangent Kernel Perspective
$f_{\mathbf{W}_{(k)}}(\mathbf{x})=f_{\mathbf{W}_{0}}(\mathbf{x})+\langle\nabla f_{\mathbf{W}_{0}}(\mathbf{x}),\mathbf{W_{(k)}}-\mathbf{W}_{0}\rangle+o(\|\mathbf{W_{(k)}}-\mathbf{W}_{0}\|_{\text{F}}^{2}),$ (12)
where $o(\|\mathbf{W_{(k)}}-\mathbf{W}_{0}\|_{\text{F}}^{2})$ is the small random quantity that tends to 00 as network width gets close to infinity, measuring the distance between updated model parameter and its initialization in Frobenius norm. Specifically, it can be shown $\|\mathbf{W_{(k)}}-\mathbf{W}_{0}\|_{\text{F}}^{2}\leq\mathcal{O}(\frac{1}{\sqrt{M}})$ with sufficiently large enough $M$. (For instance, see Remark $3.1$ in Du et al. (2018).) Under this setting, the right-hand side of (12) is linear in the network parameter $\mathbf{W}_{(k)}$. As a consequence, training on $\ell_{2}$-loss with gradient descent leads to kernel regression solution with respect to the (random) kernel induced by the feature mapping $\phi(\mathbf{x}):=\nabla_{\mathbf{W}_{0}}f(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{X}$. The inner-product of two feature mappings evaluated at two data points $\mathbf{x}_{i},\mathbf{x}_{j}$ is denoted as $\mathbf{K}^{(M)}(\mathbf{x}_{i},\mathbf{x}_{j}):=\langle\phi(\mathbf{x}_{i}),\phi(\mathbf{x}_{j})\rangle$ for all $1\leq i,j\leq n$.
Setting ReLU activation function (i.e., $\sigma(\mathbf{x})=\max(\mathbf{x},0)$), the explicit form of $\mathbf{K}^{(M)}(\mathbf{x}_{i},\mathbf{x}_{j})$ is written as follows :
$\mathbf{K}^{(M)}(\mathbf{x}_{i},\mathbf{x}_{j}):=\frac{1}{M}\sum_{r=1}^{M}a_{r}^{2}\mathbf{x}_{i}\mathbf{x}_{j}\mathbb{I}\{w_{r}^{\top}\mathbf{x}_{i}\geq 0,w_{r}^{\top}\mathbf{x}_{j}\geq 0\}+\frac{1}{M}\sum_{r=1}^{M}\mathbb{I}\{w_{r}^{\top}\mathbf{x}_{i}\geq 0,w_{r}^{\top}\mathbf{x}_{j}\geq 0\}.$
Note that the $\mathbf{K}^{(M)}(\cdot,\cdot)$ is a random matrix with respect to the initialization $a_{r}$ and $\mathbf{w}_{r}$ defined over $\mathcal{X}\times\mathcal{X}$. It is shown that it converges to its deterministic limit (i.e., $M\rightarrow{\infty}$) in probability pointwisely (Jacot et al., 2018; Arora et al., 2019b; Lee et al., 2018) and uniformly (Lai et al., 2023) over the feature space $\mathcal{X}$. In literature, the limit matrix is named NTK denoted as $\{\mathbf{K}^{\infty}(\mathbf{x}_{i},\mathbf{x}_{j})\}_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$. Hereafter, we write the eigen-decomposition of $\mathbf{K}^{\infty}=\sum_{j=1}^{n}\lambda_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{\top}$, where $\lambda_{1}\geq\dots\geq\lambda_{n}\geq 0$ with corresponding eigenvectors $\mathbf{v}_{j}\in\mathbb{R}^{n}$.
Many papers have come out in the sequel to tackle the optimization and generalization properties of neural networks in the NTK regime. Notably among them, Du et al. (2018) proves the linear convergence of training loss of shallow ReLU networks when the network parameters are updated through gradient descent with fixed step size $\eta=\mathcal{O}\big{(}\frac{\lambda_{0}}{n^{2}}\big{)}$. Specifically, the authors randomly initialized $a_{r}\sim\text{Unif}\{-1,+1\}$ and $\mathbf{w}_{r}\sim\mathcal{N}(0,\mathcal{I})$, and train the $\mathbf{w}_{r}$ fixing the output layer, leading $\mathbf{K}^{\infty}(\mathbf{x}_{i},\mathbf{x}_{j}):=\mathbb{E}_{\mathbf{W}\sim\mathcal{N}(0,\mathcal{I})}\big{[}\mathbf{x}_{i}\mathbf{x}_{j}\mathbb{I}\{w^{\top}\mathbf{x}_{i}\geq 0,w^{\top}\mathbf{x}_{j}\geq 0\}\big{]}$. Here, the linear convergence rate means that the training loss at $k$-th gradient descent update decays at a geometric rate with respect to the initial training loss, which is explicitly stated in Theorem 4.1. in Du et al. (2018) as:
Their result requires the network width $M$ to be in the order of $\Omega\big{(}\frac{n^{6}}{\lambda_{n}}\big{)}$, and the decay rate is dependent on the minimum eigenvalue of NTK, $\lambda_{n}$. Here, for the geometric decay rate, $\lambda_{n}$ needs to be strictly greater than 00.
Afterwards, there have been several attempts to reduce the overparametrization size. One of the works we are aware of is Song and Yang (2019) where the authors used matrix Chernoff bound to reduce the width size up to $M=\Omega\big{(}\frac{n^{2}}{\lambda_{n}^{4}}\big{)}$ with slightly stronger assumptions than data non-parallel assumption. Several subsequent works Allen-Zhu et al. (2018); Du et al. (2019); Zou et al. (2018); Wu et al. (2019); Oymak and Soltanolkotabi (2020); Suh et al. (2021) extended the results showing the linear convergence of training loss of deep ReLU networks with $L$-hidden layers. For a succinct comparison of the overparametrized conditions on $M$ derived in aforementioned papers, we direct readers to consult Table 1 in Zou and Gu (2019).
Motivated by the result (13), researchers further studied the spectral bias of deep neural network models, investigating why the neural dynamics learn the lower frequency components of the functions faster than they learn the higher frequency counterparts. The specific results are stated in terms of eigenvalues $\mu_{1}\geq\mu_{2}\geq\dots$ and corresponding orthonormal eigenfunctions $\phi_{1}(\cdot),\phi_{2}(\cdot),\dots$ of integral operator $\mathcal{L}_{\mathbf{K}^{\infty}}$ induced by $\mathbf{K}^{\infty}$:
$\mathcal{L}_{\mathbf{K}^{\infty}}(f)(\mathbf{x}):=\int_{\mathcal{X}}\mathbf{K}^{\infty}(\mathbf{x},\mathbf{y})f(\mathbf{y})\rho(\mathbf{dy}),\quad\forall f\in\mathcal{L}^{2}(\mathcal{X}).$
Specifically, Cao et al. (2019); Bietti and Mairal (2019) provided the spectral decay rates of $(\mu_{k})_{k}$ for shallow ReLU networks when $x\in\mathcal{S}^{d-1}$ equipped with uniform measure $\mathbf{\tau}$ (i.e., $\mathbf{x}\sim\text{Unif}(\mathcal{S}^{d-1})$). [^2] Note that eigenvalues of $\mathbf{K}^{\infty}$ (i.e.,$\{\lambda_{i}\}_{i=1}^{n}$) and eigenvalues of $\mathcal{L}_{k}$ (i.e.,$(\mu_{k})_{k}$) are different.
Proposition 3.1.
$\begin{cases}\mu_{0},\mu_{1}=\Omega(1),\\ \mu_{K}=0,\quad\text{when $k$ is odd},\\ \mu_{k}=\Omega(k^{-d-1}),\quad k\gg d.\\ \end{cases}$
The decay rate is exponentially fast in input dimension $d$. An interesting benefit of having a specific decay rate is that we can measure the size of reproducing kernel Hilbert spaces (RKHS) induced from the kernel $\mathbf{K}^{\infty}$. The slower the decay rate is, the larger the RKHS becomes, allowing higher-frequency information of function to be included in the function space. This can be easily seen through the definition of RKHS in the unit sphere:
:=$\mathcal{H}^{\infty}:=\bigg{\{}f=\sum_{k\geq 0,\mu_{k}\neq 0}\sum_{j=1}^{\mathcal{N}(d,k)}a_{k,j}Y_{k,j}(\cdot)\quad\text{ s.t }\quad\|f\|_{\mathcal{H}^{\infty}}^{2}=\sum_{k\geq 0,\mu_{k}\neq 0}\sum_{j=1}^{\mathcal{N}(d,k)}\frac{a_{k,j}^{2}}{\mu_{k}}<\infty\bigg{\}},$
where $Y_{k,j}$ for $j=1,\dots,\mathcal{N}(d,k)$ are linearly independent spherical harmonic polynomials of degree $k$ in $d$ variables. (We refer Frye and Efthimiou (2012) for more detailed expositions on spherical harmonic polynomials.) The slower the $\mu_{k}$ decays, the higher $k$-degree polynomials can be included in the $\mathcal{H}^{\infty}$.
With the specified eigendecay rates on $\mu_{k}$, Theorem $4.2$ in Cao et al. (2019) proved the spectral bias of neural network training in the NTK regime. Specifically, as long as the network is wide enough and the sample size is large enough, gradient descent first learns the target function along the eigen directions of NTK with larger eigenvalues, and learns the rest components corresponding to smaller eigenvalues:
$\frac{1}{\sqrt{n}}\left\|\mathbf{V}_{k}^{\top}(\mathbf{y}-\widehat{\mathbf{y}}^{(T)})\right\|_{2}\leq\frac{2}{\sqrt{n}}\big{(}1-\mu_{k}\big{)}^{T}\left\|\mathbf{V}_{k}^{\top}\mathbf{y}\right\|_{2},$
where $\mathbf{V}_{k}:=(n^{-1/2}\phi_{j}(\mathbf{x}_{i}))_{n\times k}$. The term $\|\mathbf{V}_{k}^{\top}(\mathbf{y}-\widehat{\mathbf{y}}^{(T)})\|_{2}$ roughly represents the speed of gradient descent learns the components of target function corresponding to the first $k$ eigenvalues. (See Lemma $4.1$ in Cao et al. (2019).) Here, we denote $\mathbf{y}:=(f^{\star}(\mathbf{x}_{1}),\dots,f^{\star}(\mathbf{x}_{n}))$ where $f^{\star}$ is a target function. Similary Hu et al. (2019) showed that gradient descent learns the linear component of target functions in the early training stage. But crucially they do not require the network to have a disproportionately large width, and the network is allowed to escape the kernel regime later in training.
Generalization of neural nets in NTK regime. Here, we review some important works that study the generalizabilities of (9). To the best of our knowledge, Arora et al. (2019a) provided the first step in understanding the role of NTK in the generalizability of neural nets. Specifically, they showed that for $M=\Omega(\frac{n^{2}\log(n)}{\lambda_{\text{min}}})$ and $k\geq\tilde{\Omega}\big{(}\frac{1}{\eta\lambda_{\text{min}}}\big{)}$, the $\ell_{2}$ population loss of $f_{\mathbf{W}_{(k)}}(\mathbf{x})$ is bounded by:
$\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim\mathcal{D}}\big{[}(f_{\mathbf{W}_{(k)}}(\mathbf{x})-\mathbf{y})^{2}\big{]}\leq\mathcal{O}\bigg{(}\sqrt{\frac{\mathbf{y}^{\top}(\mathbf{K}^{\infty})^{-1}\mathbf{y}}{n}}\bigg{)}.$ (14)
Observe the nominator in the bound can be written as $\mathbf{y}^{\top}(\mathbf{K}^{\infty})^{-1}\mathbf{y}:=\sum_{i=1}^{n}\frac{1}{\lambda_{i}}(\mathbf{v}_{i}^{\top}\mathbf{y})^{2}$. This implies the projections $\mathbf{v}_{i}^{\top}\mathbf{y}$ that correspond to small eigenvalues $\lambda_{i}$ should be small for good generalizations on unseen data. Indeed, they performed the empirical experiments on MNIST and CIFAR-10 datasets, showing the projections $\{(\mathbf{v}_{i}^{\top}\mathbf{y})\}_{i=1}^{n}$ sharply drops for true labels $\mathbf{y}$, whereas the projections are close to being uniform for random labels $\mathbf{y}$. (See Figure $1$ in their paper.)
$\mathbb{E}\big{[}\|f_{W_{(T)}}(\mathbf{x})-f^{\star}(\mathbf{x})\|_{L_{2}}^{2}\big{]}\leq\mathcal{O}\big{(}T^{-\frac{2r\beta}{2r\beta+1}}\big{)},$ (15)
where $f^{\star}\in\mathcal{L}_{\mathbf{K}^{\infty}}^{r}(\mathcal{L}_{2}(\rho_{\mathbf{x}}))$ with $r\in[1/2,1]$. The rate is minimax optimal, which is faster than $\mathcal{O}(\frac{1}{\sqrt{n}})$ in Arora et al. (2019a). The rate is characterized by two control parameters, $\beta$ and $r$, where $\beta>1$ controls the complexity of $\mathcal{H}^{\infty}$ and $r\in[1/2,1]$ controls the complexity of subset of $\mathcal{H}^{\infty}$ where the target function $f^{\star}$ belongs. Note that there exists an interesting bias-variance trade-off between these two quantities $\beta$ and $r$. For large $\beta$, the whole space $\mathcal{H}^{\infty}$ becomes small, and the subspace of $\mathcal{H}^{\infty}$ needs to be as large as possible for the faster convergence rate, and vice versa.
As noted by the following work Hu et al. (2021), the rate in (15) requires the network width to be exponential in $n$. (They didn’t specify the dependence in the main result, but can be checked in their proof.) The work Hu et al. (2021) reduced the size of overparametrization to $\tilde{\Omega}(n^{6})$ when the network parameters are estimated by gradient descent. The paper proved that the overparametrized shallow ReLU networks require $\ell_{2}$-regularization for GD to achieve minimax convergence rate $\mathcal{O}(n^{-\frac{d}{2d-1}})$. Note that this result is consistent with (15) as it is assumed $f^{\star}\in\mathcal{H}^{\infty}$ (i.e., $r=1/2$), and $\beta=\frac{d}{d-1}$ (see Lemma $3.1$ in Hu et al. (2021)). Afterward, Suh et al. (2021) extended the result to multi-layer deep ReLU networks in the NTK regime, showing that $\ell_{2}$-regularization is also required for achieving minimax rate for deep networks.
3.2 Limitations of NTK and Beyond the Kernel Regime
@@
Beyond the kernel regime. The above two works explained the superiority of neural networks over the networks in the NTK regime under some highly stylized settings. There also have been several theoretical attempts (Bai and Lee, 2019; Allen-Zhu et al., 2019) to explain how the networks estimated through gradient-based methods generalize well to the unseen data but critically do not rely on the linearization of network dynamics. The paper Bai and Lee (2019) studied the optimization and generalization of shallow networks with smooth activation function $\sigma(\cdot)$ via relating the network dynamics $f_{\mathbf{W}}(\mathbf{x})$ with higher-order approximations. The key idea is to find the subset of $\mathbf{W}:=\{(a_{r},w_{r})\}_{r=1}^{M}$ such that the dominating term in Taylor expansion (12) is not the linear term (i.e., $\langle\nabla f_{\mathbf{W}_{0}}(\mathbf{x}),\mathbf{W}-\mathbf{W}_{0}\rangle$) but a quadratic term (i.e., $\|\mathbf{W}-\mathbf{W}_{0}\|_{\text{F}}^{2}$). The authors find such $\mathbf{W}$ by running SGD on the tailored-regularized loss:
@@
The expectation in loss is taken over the diagonal matrix $\mathbf{\Sigma}$ with $\Sigma_{rr}\sim\text{Unif}\{-1,1\}$ over $r\in[M]$ and it has the effects of dampening out the linear term. The regularizer $\|\cdot\|_{2,4}^{8}$ controls the distance of moving weights to be $\mathcal{O}(M^{-1/4})$. The authors showed the landscape of optimization problem (16) has a very nice property that every second-order stationary point of the loss is nearly global optimum, despite the non-convexity of (16). These points can be easily found by the SGD algorithm. Given that $M$ is sufficiently large enough, i.e., $M\gtrsim n^{4}$, they further provided the comparisons of sample complexities between quadratic and linear models under three different scenarios. The ground truth functions estimated in the three scenarios are (1) $f^{\star}(\mathbf{x})=\alpha(\beta^{\top}\mathbf{x})^{p}$ for $\alpha\in\mathbb{R}$, $\beta\in\mathbb{R}^{d}$, and $p=1\text{ or even}$. (2) $f^{\star}(\mathbf{x})=\mathbf{x}_{1}\mathbf{x}_{2}$, and (3) $f^{\star}(\mathbf{x})=\langle\Theta^{\star},\mathbf{x}\mathbf{x}^{\top}\rangle$ where $\Theta^{\star}\in\mathbb{R}^{d\times d}$ is a rank-$r$ matrix. We summarize the results in the below table.

### v3

@@ Appendix B Classification tasks via fully-connected networks
Lastly, a paper Hu et al. (2020a) studied the convergence rate of excess risk of neural networks under the framework of Mammen and Tsybakov (1999). Specifically, the authors of the paper consider a Teacher-Student framework. In this setup, one neural network, called student net, is trained on data generated by another neural network, called teacher net. Adopting this framework can facilitate the understanding of how deep neural networks work as it provides an explicit target function with bounded complexity. Furthermore, assuming the target classifier to be a teacher network of an explicit architecture might provide insights into what specific architecture of the student classifier is needed to achieve an optimal excess risk. Under this setting, the rate of convergence is derived as $\tilde{\mathcal{O}}_{d}(n^{-\frac{2}{3}})$ for the excess risk of the empirical $0-1$ loss minimizer, given that the student network is deeper and larger than the teacher network. The authors use the notation $\tilde{\mathcal{O}}_{d}$ to denote the rate is dependent on input dimension $d$ in a logarithmic factor. When data are separable, the rate improves to $\tilde{\mathcal{O}}_{d}(n^{-1})$. In contrast, as it has already been shown in Mammen and Tsybakov (1999), under CAD assumption, the optimal rate is $\mathcal{O}(n^{-\alpha(q+1)/[\alpha(q+2)+(d-1)q]})$. Clearly, this rate suffers from “Curse of Dimensionality” but interestingly, coincides with the rate $\tilde{\mathcal{O}}_{d}(n^{-\frac{2}{3}})$ when $\alpha=1$ and $\beta\rightarrow{\infty}$. If we further allow $\alpha\rightarrow{\infty}$ (corresponding to separable data), the classical rate above recovers $\mathcal{O}(n^{-1})$. The paper also got the rate is un-improvable in a minimax sense by showing that the minimax lower bound has the same order as that of the upper bound (i.e., $\tilde{\mathcal{O}}_{d}(n^{-\frac{2}{3}})$).
Appendix C Intuition of PDE Equation (Distributional Dynamics) in Mei et al. (2018b)
In the Mean-Field setting, recall the trajectories of empirical distribution of $\theta^{(k)}$ denoted as $\widehat{\rho}^{(M)}_{k}:=\frac{1}{M}\sum_{r=1}^{M}\delta_{\theta^{(k)}_{r}}$ weakly converges to the deterministic limit $\rho_{t}\in\mathcal{P}(\mathbb{R}^{\mathcal{D}})$ as $k\rightarrow{\infty}$ and $M\rightarrow{\infty}$. The measure $\rho_{t}$ is the solution of the following nonlinear partial differential equation (PDE):
The PDE in Equation. (28) is often referred as Distributional Dynamics (DD). The intuition of the above PDE can be understood through the derivation process. To aid the understandings of readers, we only present a high-level idea for the derivation without rigor. For each particle $\theta_{r}$, we can write the $k$th update of one-pass SGD as follows:
$\theta_{r}^{k+1}=\theta_{r}^{k}+2\eta_{k}\nabla_{\theta_{r}}\sigma_{\star}(\mathbf{x}_{k},\theta_{r}^{k})\bigg{\{}\mathbf{y}_{k}-\frac{1}{M}\sum_{r=1}^{M}\sigma_{\star}(\mathbf{x}_{k};\mathbf{\theta}_{r}^{k})\bigg{\}}.$
By fixing the learning rate $\eta_{k}:=\frac{\varepsilon}{2}$ with $\varepsilon$ close to 00 and taking expectation with respect to data $(\mathbf{x},\mathbf{y})$ on both sides of the equality, we have the ODE desribing the continuum motion of each particle $\theta_{r}$:
$\dot{\mathbf{\theta}}_{r}(t)=-2\nabla_{\mathbf{\theta_{r}}}\mathcal{V}(\mathbf{\theta_{r}})-\frac{2}{M}\sum_{k=1}^{M}\nabla_{\mathbf{\theta_{r}}}\mathcal{U}(\mathbf{\theta_{r}},\mathbf{\theta_{k}}).$ (29)
Each particle $(\theta_{r})$ moves to the directions where it minimizes $-\mathcal{V}(\mathbf{\theta_{r}}):=\mathbb{E}\{\mathbf{y}\sigma_{\star}(\mathbf{x};\mathbf{\theta})\}$ to secure the best fit with the response data $\mathbf{y}$. At the same time, it moves to the directions where it minimizes $-\frac{1}{M}\sum_{k=1}^{M}\mathcal{U}(\mathbf{\theta_{r}},\mathbf{\theta_{k}})$ so that the particle $\mathbf{\theta}_{r}$ does not get too close to the other particles $(\mathbf{\theta}_{k})_{k\neq r}$. Recalling the definition of $\mathbf{\Psi}(\mathbf{\theta};\rho_{t})$, we rewrite the ODE Equation. (29) in terms of empirical distribution $\widehat{\rho}^{(M)}_{k}$:
$\dot{\mathbf{\theta}}_{r}(t):=-\nabla_{\theta_{r}}\mathbf{\Psi}(\mathbf{\theta}_{r};\widehat{\rho}^{(M)}_{k}).$ (30)
Then, the authors introduced new $M$ i.i.d. trajectories $(\bar{\theta}_{r}(t))$ with $r=1,\dots,M$ with same initializations as for the SGD (i.e., $\bar{\theta}_{r}(0)={\theta}_{r}(0)$), where the trajectory $\bar{\theta}_{r}(t)$ is the solution of the ODE $\dot{\mathbf{\bar{\theta}}}(t):=-\nabla_{\bar{\theta}}\mathbf{\Psi}(\bar{\theta}(t);\rho_{t})$, then proved $\bar{\theta}_{r}(t)$ is actually the solution of the PDE in Equation. (28). Finally, they showed the distance between two trajectories $\bar{\theta}_{r}(t)$ and $\theta_{r}(t)$ can be controlled. In other words, the above PDE Equation. (28) describes the evolution of each particle $(\theta_{r})$ in the force field created by the densities of all the other particles.
Appendix D Beyond the Kernel Regime
@@
Beyond the kernel regime. The above two works explained the superiority of neural networks over the networks in the NTK regime under some highly stylized settings. There also have been several theoretical attempts (Bai and Lee, 2019; Allen-Zhu et al., 2019) to explain how the networks estimated through gradient-based methods generalize well to the unseen data but critically do not rely on the linearization of network dynamics. The paper Bai and Lee (2019) studied the optimization and generalization of shallow networks with smooth activation function $\sigma(\cdot)$ via relating the network dynamics $f_{\mathbf{W}}(\mathbf{x})$ with higher-order approximations. The key idea is to find the subset of $\mathbf{W}:=\{(a_{r},w_{r})\}_{r=1}^{M}$ such that the dominating term in Taylor expansion Equation. (10) is not the linear term (i.e., $\langle\nabla f_{\mathbf{W}_{0}}(\mathbf{x}),\mathbf{W}-\mathbf{W}_{0}\rangle$) but a quadratic term (i.e., $\|\mathbf{W}-\mathbf{W}_{0}\|_{\text{F}}^{2}$). The authors find such $\mathbf{W}$ by running SGD on the tailored-regularized loss:
@@
The expectation in loss is taken over the diagonal matrix $\mathbf{\Sigma}$ with $\Sigma_{rr}\sim\text{Unif}\{-1,1\}$ over $r\in[M]$ and it has the effects of dampening out the linear term. The regularizer $\|\cdot\|_{2,4}^{8}$ controls the distance of moving weights to be $\mathcal{O}(M^{-1/4})$. The authors showed the landscape of optimization problem Equation. (31) has a very nice property that every second-order stationary point of the loss is nearly global optimum, despite the non-convexity of Equation. (31). These points can be easily found by the SGD algorithm. Given that $M$ is sufficiently large enough, i.e., $M\gtrsim n^{4}$, they further provided the comparisons of sample complexities between quadratic and linear models under three different scenarios. The ground truth functions estimated in the three scenarios are (1) $f^{\star}(\mathbf{x})=\alpha(\beta^{\top}\mathbf{x})^{p}$ for $\alpha\in\mathbb{R}$, $\beta\in\mathbb{R}^{d}$, and $p=1\text{ or even}$. (2) $f^{\star}(\mathbf{x})=\mathbf{x}_{1}\mathbf{x}_{2}$, and (3) $f^{\star}(\mathbf{x})=\langle\Theta^{\star},\mathbf{x}\mathbf{x}^{\top}\rangle$ where $\Theta^{\star}\in\mathbb{R}^{d\times d}$ is a rank-$r$ matrix. We summarize the results in the below table.